Train Travel. A train leaves Barstow traveling east at 35km/h. One hour later, a faster train leaves Barstow, also traveling east on a parallel track at 40km/h. How far from Barstow will the faster train catch up with the slower one?

Answer by Alan3354(21605)

A train leaves Barstow traveling east at 35km/h. One hour later, a faster train leaves Barstow, also traveling east on a parallel track at 40km/h. How far from Barstow will the faster train catch up with the slower one?
------------------------
The 1st train is 35 km away when the 2nd train leaves.
The 2nd train gains on the slower train at 5 kph.
35/5 = 7 hours to overtake it.
In 7 hours the 2nd train will be 40*7 = 280 km from Barstow.
---------------
As a check, the slower train will be 35*8 = 280 km from Barstow also.